Research & Science Home ESA Public Web Site Sci-Tech Portal Herschel Public Web Site Herschel Sci-Tech Portal
Astrophysics Missions Planetary Exploration Missions Solar Terrestrial Science Missions Fundamental Physics Missions Science Faculty
Divisions
Herschel General Information
Herschel Science Centre Home
Latest News
Mission Overview
Science Instruments
Conferences/Workshops
News & Press Releases
e-News
Herschel Science Team
Herschel Users' Group
Herschel People
Useful links
Herschel Announcement of Opportunity (AO-2)
Introduction
'How-to' step-by-step
Documentation
Tools
AO Latest News
Herschel Observing
Overview
Calls for proposals
Scheduling
Observing status
Proprietary time
Observing Log
Key Programmes
AO-1 Programmes (GT1 + OT1)
AO-2 Programmes (GT2 + OT2)
"Must-Do" Programmes
Filler Programmes
Herschel Documentation
Observatory
HIFI
PACS
SPIRE
Herschel Data Products
Data Products Overview
Known Issues
Herschel Science Archive (HSA)
User Provided Data Products
Postcard Gallery - browse HSA
Herschel Data Processing
Data Processing Overview
HIPE Download
User Contributed Software
HIPE Community
HIPE Video Tutorials
Herschel Publications
Publishing Rules & Guidelines
Scientific Publications
Communicating Your Results
Herschel User Services
Services Overview
Helpdesk
Proposal Handling
Subscribe to Herschel eMail list
Herschel User Registration
User Registration
Lost Password ???

HSC Operations (B)Log

DISCLAIMER: This blog is since 2009 no longer active.

Welcome to the Herschel Science Centre Operations (B)Log. Here members of the HSC will provide frequent updates to keep you informed and for you to share in the excitement as the mission progresses.

A note about dates: The Herschel launch date was 14 May 2009, this is day L, this was also Day Of Year (DOY) 134, we provide all since various sources use different time references. There is also another concept, the Operational Day (OD), which is not a calendar day but has a duration of approximately 24 hours.


25 November 2009 (L+195, DOY 329)

The findings of the HIFI investigation team and the proposed plan for bringing HIFI back in operation were presented to the ESA Director of Science and Robotic Exploration and others in a briefing meeting held in ESTEC on 25 November 2009. The plan presented was endorsed, and is now the adopted way forward.

[G. Pilbratt, Herschel Project Scientist, posted 27 November 2009]


19 November 2009 (L+189, DOY 323)

The plan for bringing HIFI back in operation is now on the table, and subject to formal approval in the coming week. The plan consists of performing a number of software updates in the coming weeks, followed by a full instrument switch-on in January 2010, after which HIFI operations will resume. See information provided on the SRON website.

[G. Pilbratt, Herschel Project Scientist, posted 20 November 2009]


26 October 2009 (L+165, DOY 299)

The HIFI failure investigation team set-up at ESA in support of the Principal Investigator effort in SRON is finalising its work and a draft report is currently under review within the team. The content of this report will be presented to the ESA Director General and the Director of Science and Robotic Exploration this week. Due to the capability to perform extended tests on fully representative hardware and software on the ground, and to the excellent cooperation between the PI team and ESA specialists, the investigation has arrived at a complete and consistent failure scenario which can explain the observable evidence. Some further consolidating investigations on this scenario are still on-going but the picture that has emerged is the following:

A Single Event Upset (SEU) in the Random Access Memory (RAM) of the Local Oscillator Control Unit (LCU) microcontroller would be at the origin of a chain reaction that eventually results in an unplanned emergency switch off of sensitive instrument components. Designed to protect the local oscillators against damage from undervoltage on the spacecraft 28 V bus, this switch was thrown while the 28 V bus was alive. The resulting load transient in the internal power system of the LCU associated with this mode change created a stress in the power converters, leading to a permanent failure of one of the diodes.

Before the restart of the instrument on the redundant signal chain can be performed, changes in the operation and protection logic of the instrument must be implemented and validated in on-board software to prevent re-occurrence of such a sequence of events.

Although the detailed investigations have uncovered some marginalities in the stress applied to certain diodes in the internal power system of the LCU, laboratory tests have shown these diodes to be quite resilient against the short overvoltage spikes they are subjected to during nominal instrument operations. The investigation team is confident that HIFI can perform nominally for the remainder of the mission if the required corrective actions are implemented.

[J. Riedinger, Herschel Mission Manager and co-chair of the ESA investigation team]


5 October 2009 (L+144, DOY 278)

Progress towards re-enabling HIFI

Today, SRON have added an entry to their top level web page (www.sron.nl), in which they report about the progress that is being made towards re-enabling HIFI operations. Despite the progress which is undoubtedly made by the joint HIFI/ESA investigation team, the HIFI Project Manager cautions that switching the instrument back on may still be a few weeks in the future: We must ensure that we have taken all conceivable measures to minimise the risk of such a chain of events from happening again. One of these measures may require changes in on-board software, the validation of which would have to be performed with the utmost care.

[J. Riedinger, Herschel Mission Manager]


29 September 2009 (L+138, DOY 272)

First science data distributed to the users community

Yesterday 28 September Herschel reached another important milestone: the first set of observations corresponding to the so-called 'Science Demonstration Phase" were made available to their owners. This delivery took place several weeks ahead of the originally planned schedule and marks the start of the transition from the Performance Verification Phase to the Routine Phase of operations.

The observations (six so far only) were conducted on 12 September 2009 and successfully processed with the latest version of the pipeline as implemented in a dedicated version of HIPE that was also distributed to the observers. They are SPIRE scan maps of a variety of astronomical sources including a galactic HII region, a proto-planetary nebula, a supernova remnant, and a couple of galaxy clusters. This is the first of the observing modes released to the users community following its early validation during the Performance Verification Phase.

If there are no major contingencies, we expect to be running the Science Demonstration Phase at full speed during the second half of October and the whole month of November, as initially planned, first with PACS and SPIRE AORs, and then with HIFI as well, once the instrument resumes operations and completes its own performance and verification phase, delayed because of the problems already reported in this (B)log.

[P. Garcia-Lario, HSC ESAC, posted 30 September 2009]


11 September 2009 (L+120, DOY 254)

Update on the HIFI anomaly investigation

On Monday this week, the team established by ESA to investigate into the HIFI anomaly together with SRON met with instrument experts in Groningen for an extended question and answer session. Over the coming weeks the ESA team will help HIFI in trying to piece together what appears to be a complex puzzle of events which includes, but seems to be more complex than, failure of a DC/DC converter. By correctly understanding the failure scenario we hope to be able to minimise the risk when we switch HIFI back on, possibly using its redundant signal chain.

[J. Riedinger, Mission Manager and co-chair of the ESA investigation team]


4 September 2009 (L+113, DOY 247)

An update regarding the ongoing HIFI LCU malfunction investigation activities has today been posted on the ESA Corporate website.

The bottom line is that the investigation effort underway in SRON since the problem occurred is being augmented by an ESA team. The common objective is to determine the chain of events leading to the malfunction, and to decide when and under what conditions it can be considered safe to continue HIFI operations using the redundant set of warm electronics units.

[G. Pilbratt from ESTEC, posted 4 September 2009]


14 August 2009 (L+92, DOY 226)

Statement regarding HIFI LCU malfunction

On 3 August 2009 it was discovered by HIFI that on the day before the local oscillator control unit (LCU) had developed an anomaly leading to HIFI shutting down. Since then the HIFI consortium has been intensely investigating the nature of the problem; today the HIFI Principal Investigator Frank Helmich has provided the following statement:

"During normal check-out observations HIFI was shut down due to an unknown event in the analog voltage supply unit of the control electronics of the HIFI Local Oscillator. Engineers and scientists from the HIFI Instrument Control Centre and the hardware institutes are working to determine the cause and resume HIFI observations. An ongoing analysis suggests that there was either an upset in the power supply unit itself, or a voltage drop of the central unit supplying power to it. The power supply unit is needed for the instrument to obtain science observations. HIFI does have a duplicate, or redundant, set of electronics that will be used and which provides full functionality. The engineering and science team is working to understand the cause of the anomaly before switching to the duplicate set of electronics."

[G. Pilbratt from IAU GA in Rio de Janeiro, posted 14 August 2009]


10 July 2009 (L+57, DOY 191)

Herschel 'First Light' observations

You can now have a look at the 'first-light' observations released today, including SPIRE imaging of the nearby galaxies M66 and M74, HIFI spectroscopy of the star-forming region DR21, and PACS imaging spectroscopy of the planetary nebula NGC 6543.

The beauty and quality of these very first test observations is encouraging and demonstrates that a lot of new discoveries and exciting science is ahead of us! Follow Herschel progress also on the 'Latest News' page.

[P. Garcia-Lario, HSC ESAC, posted 10 July 2009]


7 July 2009 (L+54, DOY 188)

'First Light' web release this friday

A collection of 'First Light' images and spectra taken by the three instruments onboard Herschel during the days following the cryo-cover opening will be publicly made available this Friday 10 July. Watch out for the accompanying set of coordinated web releases by ESA and the different ICC consortia!

[P. Garcia-Lario, HSC ESAC, posted 7 July 2009]


3 July 2009 (L+50, DOY 184)

The coldest place in the outer space is no longer Herschel

Last night, the detectors of Planck's HFI (High Frequency Instrument) reached a temperature of 100 mK, making them the coldest thing in space, surpassing the record established by SPIRE detectors, which are operating at a warmer temperature of only 280 mK. After a successful orbit insertion maneouvre started at 13:15 CEST yesterday, Planck has now entered its final orbit around L2. Congratulations to our Planck colleagues!

[P. Garcia-Lario, HSC ESAC, posted 6 July 2009]


2 July 2009 (L+49, DOY 183)

50 days in space

Today is the 50th day of Herschel in space. As of yesterday, 80% of the planned commissioning phase activities have been executed. The initial check-out comprises a large number of tests, including switching-on of all instruments, basic functional tests, controlled cooling of the telescope, local oscillator stability measurements by HIFI, determination of the cooler recycling hold times by PACS and SPIRE, cryo-cover opening and initial determination of the focal plane geometry following 'sneak previews' of the infrared sky by all instruments, among many other checking activities. After a (so far) flawless period of almost 2 months of Commissioning Phase, with Herschel functioning nominally, we are now looking forward to starting the Performance Verification phase activities formally, something currently expected to occur on 16 July.

[P. Garcia-Lario, HSC ESAC, posted 6 July 2009]


29 June 2009 (L+46, DOY 180)

The Uplink validation chain

Before any observations can be uplinked to Herschel they must go through a series of rigorous tests. The HSC is the "meat in the sandwich" between ICC and MOC when carrying out uplink validation. The aim is to ensure that Herschel is protected from receiving any commands that could, potentially, be harmful and that nothing that has not been fully validated gets even as far as MOC, let alone to the spacecraft.

The Instrument Control Centres (ICCs) usually have to prepare their observation requests under great stress and time pressure, so it is essential to ensure that everything that has been delivered is correct and self-consistent and using the latest version of the uplink software, orbit file, etc. Under pressure, it is easy to make a mistake and deliver a wrong file in the middle of dozens of correct ones. Similarly, there is information that a particular ICC does not and cannot know when it is preparing its own observations, especially when a particular day is shared between two or more instruments: how much time is needed to slew from the end position of the first instrument's observations to the start position of the other's? Is the slew pattern acceptable to Flight Dynamics? Do all the observations from the different instruments fit in to the time available when combined?

If an inconsistency is found in validation, it must be understood and, almost invariably, the delivery is rejected and must be re-delivered. Only when the uplink information has been fully validated does it then go to the HSC Mission Planners who put together the final observing schedule and apply the last layer of checks, comparing between the schedule that is expected from the ICC input and the one that is actually obtained when that input is processed at the HSC. If a discrepancy is found, however trivial, it must be queried and if a change has had to be made to the sequence of observations, approval for the change must be sought from the ICC. Only when the "i"s are dotted, the "t"s crossed and everything is shipshape and Bristol fashion will the planning files be passed to the Project Scientist for his approval of the schedule and then on to MOC for transmission to the spacecraft.

The HSC Uplink Validation Team:
Delivery validators - Larry O'Rourke and Mark Kidger
Mission Planning - Charo Lorente and the Mission Planning Gang (Álvaro Llorente, Fernando Rodríguez and Mar Sierra)

[M. Kidger, HSC ESAC posted 6 July 2009]


25 June 2009 (L+42, DOY 176)

Can Herschel see Planck?

Some people may think that Herschel and Planck are quite close together in their orbits around the Second Lagrange Point of the Sun-Earth system and that a hypothetical observer on Herschel would be able to see Planck and vice-versa. How true is this?

At launch, of course, Herschel and Planck were very close together, but they separated quite quickly. At midnight on launch day the two were just 44 km apart and to a hypothetical stowaway on Herschel, Planck would have been about magnitude -3, but this distance between them increased quickly and, as a result, Planck would have faded rapidly from sight. By early on the morning of 17 May it would have been lost to the naked eye. On 29 May the separation between the two passed 50 000 km and on 7 June it passed 100 000 km. At the same time, the relative velocity between Herschel and Planck has increased considerably from the initial 1.3 m/s at 00UT on 15 May, to a maximum velocity of recession of 130 m/s on 18 June. In contrast, on 22 August, Planck will be approaching Herschel at no less than 119 m/s.

The distance and relative velocity between Herschel and Planck follow an approximately 3 month cycle. However, they are never separated by less than 0.5 Lunar Distances (LD) and the separation reaches a maximum value of 1.225 LD (471 000 km). Over the course of 2009, the closest approaches are on 5 September (207 000 km) and 4 December (159 000 km), with maximum separations on 26 July (471 000km) and 16 October (458 000km). Seen from Herschel, Planck is never brighter than V=14.9 and gets as faint as V=17.2

A Herschel or Planck ephemeris can be generated for any observing site on Earth or in space using JPL's Horizons system.

[M. Kidger, HSC ESAC, posted 6 July 2009]


19 June 2009 (L+36, DOY 170)

Herschel opened its infrared eyes

On 14 June 2009 PACS obtained images of M51, the 'Whirlpool Galaxy' in its first test observation immediately following cryo-cover opening. The images, obtained at 70, 100 and 160 microns are shown below and demonstrate that the optical performance of Herschel and its large telescope is meeting their high expectations.

[Image: Herschel/PACS at 160, 100, and 70 µm]

[P. Garcia-Lario, HSC ESAC, posted 6 July 2009]


14 June 2009 (L+31, DOY 165)

At 10:53:17 UTC this morning the Herschel cryocover was commanded to open by manual commanding from the MOC at ESOC.

The command was executed nominally, there is telemetry indicating that the cover reached the open position, the shaking of the spacecraft caused by the opening was seen on the gyros and the phase separator, and the temperatures at L0 and L1 slightly changed.

All of this is consistent with what you would expect from a successful cryocover opening!

[G. Pilbratt, from MOC ESOC, posted 14 June 2009]


11 June 2009 (L+28, DOY 162)

Opening of the cryocover

The opening of the cryocover is planned for Sunday 14 June. By then the telescope will have reached a temperature of about 120 K. The opening will enable the instruments inside the cryostat for the first time to see out to the real sky. You can get a very approximate idea of what is going to happen by watching the video below, available in YouTube, where the opening of the cryostat lid is shown in slow motion during one of the cryogenic qualification tests carried out by Austrian Aerospace and EADS Astrium at Friedrichshafen, Germany.

[P. Garcia-Lario from HSC ESAC, posted 11 June 2009]


10 June 2009 (L+27, DOY 161)

Herschel! Herschel! Wherefore art thou Herschel?

Since launch Herschel has made various trajectory correction manoeuvres, the last of them on 10 June and, of course, now has an accurately measured orbit, rather than the predicted orbit that had to be used initially. So, where exactly is Herschel now and where will it be in the over the next few months?

Herschel is now effectively at the 2nd Lagrange Point of the Sun-Earth system and slowing rapidly as it reaches the apogee of its orbit. At 00UT on 12 June it will be receding at 168m/s. At 00UT on 22 June at 95m/s. And at 00UT on 2 July, at just 40m/s. Eventually, at 09:30UT approximately on 7 July, Herschel will reach apogee at 1.576 million kilometres from Earth and will then start to loop backwards until 5 September when it will be at perigee at a distance of 1.248 million kilometres.

As it comes in towards perigee, Herschel will drop below the ecliptic, crossing from Ophiuchus to Serpens, then back again briefly into Ophiuchus, before spending the second half of July and first half of August in Sagittarius. Finally, two and a half weeks before perigee, it will reach a minimum declination of -30 degrees in the southern constellation of Microscopium. By perigee, Herschel will have climbed through Piscis Australis to reach declination -21 degrees in Aquarius.

Herschel's magnitude will vary between V=19.5 at apogee and 19.0 at perigee. During the summer, as it passes through the dense star clouds close to the Galactic Centre, Herschel will be an extremely difficult object to observe from Earth.

Daily updates of Herschel's position are given in the Herschel Twitter Page (ESAHerschel)

A Herschel ephemeris can be generated for any observing site on Earth or in space using JPL's Horizons system and setting the target as "Herschel Space Observatory".

[M. Kidger from HSC ESAC, posted 11 June 2009]


9 June 2009 (L+26, DOY 160)

Big helium venting nozzles closed

Yesterday 8 June the big helium venting nozzles were closed by telecommand from ESOC. Successful closure of the valves was immediately confirmed by an increase of the temperatures of the phase separator, the level 1 and the external vent line as well as the pressure in the external vent line. The cryostat behaviour is nominal.

This was the last planned activity on the helium system, from now on the cryostat will be in passive control.

[P. Garcia-Lario from HSC ESAC, posted 11 June 2009]


7 June 2009 (L+24, DOY 158)

Space weather and SREM data during Days 1-21

Herschel has launched during what has been an unprecedentedly low minimum of solar activity in modern times, with more than 150 consecutive spotless days preceding launch. It is expected that Herschel operations will cover the rise to and peak of the next solar maximum, so monitoring space weather during the mission is important, as is the level of activity at maximum; predictions have stated that the later the rise to maximum starts, the lower the maximum is likely to be.

Herschel carries the SREM (Space Radiation Environment Monitor) as a passenger (left image below), to measure constantly the radiation environment to which Herschel is exposed. Herschel Science Centre scientists analyse this data daily and are looking for possible correlations with solar activity and with predictions of future activity.

On Day 18 a sunspot group with 13 spots formed, leading to a small (B1-class) flare at 08:07UT on June 1st (OD-18), a second B1-class flare at 06:39UT on June 2nd (OD-19). These were the first two flares registered during the mission. Later, three B-class flares were seen on June 3rd (OD20-21).

SREM data (middle image below) shows that there has been no important activity registered in the SREM data since Herschel crossed the Van Allen Radiation Belts. The variations in count rate can all be attributed to noise.

The peak count rate during the Van Allen Belt crossing was about 3 orders of magnitude greater than the count rate at any other point during the transfer orbit. A closer look at the data on an expanded scale (right image below) shows no evidence of an increase in the SREM counts at the time of first two flares, which are marked with arrows.

Space Radiation Environment Monitor (SREM) mounted on Herschel SREM count rates (OD-0001 to OD-0021) SREM count rates (closer look)

[M. Kidger & M. Sánchez Portal from HSC ESAC, posted 9 June 2009]


6 June 2009 (L+23, DOY 157)

Getting ready for cryocover opening, telescope 'decontamination heating' terminated.

As a precaution against possible contamination by out-gassing from the spacecraft depositing onto the telescope during the early in-flight mission, it was decided to prevent the telescope from getting colder than 170 K by heating it. This is often referred to as 'decontamination heating'. After launch the telescope cooled down as predicted, the secondary mirror (M2) more rapidly than the primary mirror (M1), since M2 is more 'exposed' to space than M1. Consequently, the heating of M2 commenced on 18 May, 4 days after launch, followed by M1 two days later. The duration of the heating is maximized within the constraint that at the time of the cryocover opening the telescope should have cooled down to about 120 K. Based on predictions for the final telescope cooldown the decontamination heating was terminated on 6 June, 23 days after launch, in preparation for the cryocover opening which will take place on 14 June.

[G. Pilbratt from ESTEC, posted 10 June 2009]


4 June 2009 (L+21, DOY 155)

Attitude pointing responsibility transferred to the HSC

Once the spacecraft's Attitude Control & Measurement System (ACMS) has been declared 'commissioned' by the Mission Operations Centre (MOC), approval has been given to HSC to start using attitude request commands as of the start of the Daily Telecommunications Period (DTCP) of Operational Day 21, at 10:19z on 3 June. This means that from now until the end of the mission, the Mission Planning Team at the HSC is taking over attitude pointing responsibility.

[P. Garcia-Lario from HSC ESAC, posted 8 June 2009]
3 June 2009 (L+20, DOY 154)

PACS grating and chopping initial functional tests completed

The commissioning of the chopper and grating (tuning of control parameters) has been completed today with an achieved performance equal or better than what was reached on ground.

The chopper behaviour was found to be identical to what was experienced during the ground tests, giving confidence that the chosen control parameters are robust against variations in the operational temperature of the chopper.

The very good stability of the grating was also confirmed with several sets of parameters.

As of today, the manual commanding of the PACS mechanics is finalised and the two mechanisms can be declared commissioned.

[B. Altieri from MOC ESOC & R. Vavrek from PACS ICC, posted 5 June 2009]


2 June 2009 (L+19, DOY 153)

First results from SPIRE spectrometer

The picture below shows the first SPIRE interferogram, observed on May 29 using the SPIRE Fourier Transform Spectrometer. This is a result of the first successful tests of the SPIRE spectrometer in space (although currently only viewing the cryostat lid and not yet the sky!).

First SPIRE test interferogram, observed on May 29 using the SPIRE Fourier Transform Spectrometer

[SPIRE team, posted 8 June 2009]


27 May 2009 (L+13, DOY 147)

Following the successful switch on of HIFI and PACS during OD 11, all 56 PACS observations and 9 HIFI observations containing engineering data were processed successfully by the Herschel Data Processing pipeline to the expected level. Together with 6 SPIRE observations also executed during OD 11, and the available auxiliary files, they were ingested into the Herschel Science Archive, and are now available to the Herschel calibration scientists. Using for the first time in operations, the Herschel Data Processing software of the three instruments and the Herschel Science Archive system achieved a flawless end-to-end processing and archiving. Herschel has its archive populated before first light!

[S. Ott from ESAC, posted 27 May 2009]


26 May 2009 (L+12, DOY 146)

On 24 May close to 17:00 UT the third and final Herschel science instrument, HIFI, was successfully switched on. In the following 30 hours it has undergone a series of tests that show the functionality of its focal plane units -- the 4 spectrometers, its electronics units and the main moving component its chopper. All items have been shown to be working in similar fashion to on-ground testing and all functional tests have been declared successful.

As HIFI is a heterodyne unit, it also requires the use of a local oscillator unit. HIFI has a sophisticated and complex local oscillator system that allows its spectrometers to obtain high-resolution spectra across a very wide range of frequencies. The local oscillator (LO) unit, with its 14 separate signal-carrying chains,was tested today (26 May at ~14:15 - 17:00 UT). All 14 LO channels are alive and show power levels at the detectors (mixers) consistent with previous on-ground measurements. No icing of the LO windows is apparent, as was considered a possibility from outgassing during launch, and no misalignment of the LO chains is apparent to within 1-2%.

The HIFI instrument is fully functional and will now start to be setup to work optimally for science operations.

[T. Marston from MOC ESOC, posted 26 May 2009]


24 May 2009 (L+10, DOY 144)

PACS switch-on and initial functional tests

PACS has successfully been switched-on and passed the first activity in orbit, the so-called Short Functional Tests, on 24 May 2009. Results confirm that all mechanisms are moving on a very similar way than during the last ground test in Kourou, calibration sources are heated up, the sorption cooler is electrically functional. During a period of one hour the spectrometer's photoconductor arrays were fully operated, signal in the blue array is nominal, for the red array the spacecraft temperature conditions were still a bit high. Functional tests of the photometer's bolometer detectors have proven that we have no new "dead" pixels in any of the arrays.

In the coming days the bolometer's sorption cooler will be recycled before the acquisition of its first real optical signal and both spectrometer and photometer array settings will be fine tuned for in-orbit environment.

The PACS PI Albrecht Poglitsch reports "under the given boundary conditions of this phase of the mission and the scope of this test, our general impression is that things went as well as they possibly could have. Let's hope it will continue like this!"

[R.Vavrek from MOC ESOC, posted 25 May 2009]


22 May 2009 (L+8, DOY 142)

Cool news! The first SPIRE helium-3 cooler recycling has been successfully performed today. The SPIRE PI Matt Griffin reports on behalf of the SPIRE consortium that the cooler was recycled manually with an as yet unoptimised procedure. Nevertheless a temperature of 293 mK was reached, making SPIRE - at least temporarily for now - the coldest known object in outer space!

[G. Pilbratt from ESTEC, posted 22 May 2009]


21 May 2009 (L+7, DOY 141)

Following the switch-on of the SPIRE instrument, the first Herschel in-flight instrument data, 29 SPIRE Manual Commanding observations, were successfully processed by the Herschel Data Processing pipelines and ingested into the Herschel Science Archive. Achieving a flawless end to end processing and archiving with engineering observations, some of the most difficult observations to handle, is a great start for the Herschel ground segment.

[S. Ott from ESAC, posted 22 May 2009]


20-21 May 2009 (L+6-L+7, DOY 140-141)

Where is Herschel?

Herschel is moving almost directly away from the Sun towards L2 and thus nearly directly away from us, so it does not move rapidly across the sky, even though it is still moving quite quickly through space. At midnight Universal Time (GMT) on launch day it was still going at an impressive 2.29km/s, having separated at 9.7km/s. Tonight, its speed has dropped to a more sedate 0.70km/s. At 699 000km distance signals now take 2.3 seconds to reach the satellite from Earth.

Based on the predicted orbit, Herschel will reach its greatest distance from Earth on the morning of July 7th, when it will be at a distance of 2.28 million kilometers. From then on it will start to get closer again until September 5th when it will have closed to a mere 1.25 million kilometres, finally settling into an orbit about L2 with a 6-month period.

Tonight Herschel is close to Right Ascension 16h17m, Declination -02° 35' in the constellation of Serpens Caput, about 3.5 degrees south of the magnitude 4.8 star Sigma Serpentis.

However, if you want to find Herschel, it is getting quite faint now, as it is getting close to magnitude 17. You can still see it though quite easily with a 20-cm telescope and a CCD, as the image below shows.

Herschel, Planck and the Sylda observed from earth!

This image shows Herschel, Planck and the Sylda, observed from his back yard observatory in Lanzarote by the Argentinian amateur astronomer Gustavo Muler on the night of May 18/19th, with a 30-cm telescope + CCD. The telescope tracked on the satellites, so you see the stars as long trails crossing the field of view. This image was composed of 80 exposures of 30 seconds each.

You can calculate an ephemeris for Herschel for your location in the JPL Horizons system.
Enter "Herschel Space Observatory" as the target name.

[M. Kidger from HSC ESAC, posted 20 May 2009]


19-20 May 2009 (L+5-L+6, DOY 139-140)

Herschel Switches On the First Science Instrument

Over the course of the last two days, the first instrument on Herschel has been successfully switched on. This is the SPIRE instrument, which has both a photometer and a FTS spectrometer. The full arrays were able to be switched on in both cases and all looked exactly as on the ground -- and possibly a bit better. No pixels in any of the arrays were lost. The original intention was to not attempt a full turn-on of the spectrometer, but the temperature conditions were suitable and the SPIRE team took advantage of this to check out the spectrometer at an early stage. Also tested were SPIRE's beam steering mirror, used for dithering or jiggling targets on the bolometer arrays, and the FTS spectrometer mechanism. All tests were successful.

SPIRE has further testing over the next few days before the other two instruments, PACS and HIFI are switched on -- on Sunday night.
Congratulations to SPIRE on their success so far!

[T. Marston from MOC ESOC, posted 21 May 2009]


18 May 2009 (L+4, DOY 138)

The Herschel telescope decontamination heating is underway. It has been decided to prevent the Herschel telescope from cooling down to below 170 K for as long as possible compatible with opening the cryocover at a temperature of 120 K, where the cryocover opening date is determined by the needs of the science instruments. The secondary mirror (M2) is more exposed and cools more rapidly than the primary mirror (M1). Thus, having been tested on 17 May, at 06:13 (UT) on 18 May the M2 decontamination heating was started, while later the same day the M1 heating was briefly functionally verified (at the time the M1 temperature was about 187 K). The M1 heating will commence in the early hours of 20 May. The telescope heating will be discontinued about 6 days before cryocover opening to allow telescope cooldown.

A second, smaller, course correction manoeuvre was successfully performed on 18 May, with a target delta-V of 0.99 m/s.

[G. Pilbratt from ESTEC, posted 19 May 2009]


16-17 May 2009 (L+2-L+3, DOY 136-137)

Today I had occasion to record in the HSCOM log:

HSC essentially nominal. Quietest day for almost a year, as we wait for spacecraft checks to give way to more instrument activities. 2 Spacecraft (Herschel & Planck) crossing Lunar orbit. In the past 24 hours our Spitzer colleagues reported the exhaustion of their cryogen and we congratulated them on the end of their so-successful cold mission. Quite a special day!

[L. Metcalfe from HSC ESAC, posted 17 May 2009]


15-16 May 2009 (L+1-L+2, DOY 135-136)

Herschel operations are essentially nominal and the spacecraft continues to perform well. The first trajectory course correction was successfully carried out as planned about a day after liftoff on 15 May, injecting 8.7 m/s being within 1-2% of the planned magnitude, a small delta will be performed in a few days. On 16 May the spacecraft attitude control mode was successfully changed from thruster controlled mode to reaction wheel controlled mode.

The cooldown of the spacecraft follows the predictions. In particular this has been verified for the telescope which will be heated to stabilise its temperature at 170 K for an extended period (the heating will start on 18 May), and the cryostat which will reach maximum superfluid helium temperature of just under 2 K on 18 May and maximum mass flow of just under 15 mg/s on 19 May.

[G. Pilbratt from CSG Kourou, posted 17 May 2009]


14 May 2009 (Launch date, DOY 134)

Herschel was launched together with Planck on flight V188 on an Ariane 5 ECA on 14 May 2009 from CSG Kourou, the liftoff took place at 13:12:02 UT, at the very beginning of the launch window. The launch itself was flawless, and Herschel separated from the launcher at 13:37:55 UT, just under 26 mins after lift-off.

The spacecraft was acquired by the New Norcia ground station at 13:49 UT, and good telemetry was received shortly afterwards. The spacecraft was confirmed to have acquired nominal attitude and overall status was confirmed nominal. The uplink was established and the first command was executed at 14:10 UT.

Already before even separating from the launcher, a major milestone had been achieved. The helium subsystem is working according to the book, both the large and small nozzles are open, and there is positive confirmation that the phase separator is working nominally. Together with a helium temperature at launch of only 1.81 K, this is good news regarding the cryo subsystem.

The launcher performance was very good, and the preliminary assessment of the orbit is: perigee 270.0 km (intended 270.0 ± 4.5), apogee 1,197,080 km (intended 1,193,622 ± 151,800), inclination 5.99 deg (intended 6.00 ± 0.06); or in plain english: spot on.

[G. Pilbratt from CSG Kourou, posted 15 May 2009]

Last updated: Tuesday, 01-Apr-2014 11:25:30 CEST
Copyright 2000 - © European Space Agency. All rights reserved. Contact the HSC webmaster